# Green Finance and Carbon Emissions: A Nonlinear and Interaction Analysis Using Bayesian Additive Regression Trees

### v1

@@ 3.1.3 Other independent variables
In addition, carbon emissions are also influenced by a range of socioeconomic and environmental factors. Building on the existing literature from section 2, this study incorporates several variables into the model to better capture the potential relationships among key predictors. These include level of economic development, industrial structure, research and development (R&D) intensity, urbanization level, population density, foreign direct investment, degree of government intervention, environmental regulation intensity. In light of empirical studies that have confirmed the significant regional heterogeneity in the impact of green finance on carbon emissions, this study further divides the provinces into eastern, central, and western regions based on the traditional regional classification method, and incorporates this classification as a categorical variable in the model.
@@ 3.2 Data sources
This study constructs a panel dataset covering 30 provinces in China from 2000 to 2022. Due to severe data limitations, the Tibet, Hong Kong, Marco and Taiwan is excluded from the analysis. A more detailed description of the data is shown in Table 2.
@@ 3.3 Data preprocessing
To reduce right-skewness in the distribution of variables, we applied logarithmic transformations to the dependent variable (CEI) and to selected independent variables, namely economic development (GDP), population density (population), industrial structure (industry), and environmental regulation (environment), all of which exhibited skewness statistics greater than 2. We then employed Generalized Variance Inflation Factors (GVIF) to conduct a preliminary analysis of potential multicollinearity among explanatory variables. A GVIF value close to 1 indicates little to no multicollinearity, while values equal to or greater than 5 or 10 suggest strong multicollinearity obrien2007caution . Similarly, the adjusted generalized standard error inflation factor (aGSIF) values exceeding 2.2 or 3.2 are also considered indicative of significant multicollinearity (Fannie et al., 2025).
@@ 4.1 Model specification
Originally proposed by Chipman et al. chipman2010bart , Bayesian Additive Regression Trees (BART) is a flexible, nonparametric Bayesian ensemble method that represents the response variable $y$ as the sum of $m$ regression trees:
@@
The update of the tree structure $T_{i}$ is conducted using a Metropolis–Hastings (MH) step. New tree proposals are generated through operations such as growing or pruning nodes, modifying split rules, or swapping decision nodes. The acceptance probability is computed by integrating out the terminal node parameters $M_{i}$, which is tractable due to the use of conjugate priors. Once the new tree is accepted, $M_{i}$ is updated by drawing from its full conditional posterior distribution, which follows a normal distribution. The noise variance $\sigma^{2}$ is updated at the end of each iteration by drawing from a full conditional inverse-gamma distribution. By running the MCMC algorithm for a sufficiently large number of iterations and discarding the initial samples as burn-in, a sequence of fitted functions is obtained:
@@
The model is evaluated using Root Mean Square Error (RMSE) and pseudo-$R^{2}$. A lower RMSE indicates a better model fit. The RMSE is calculated as follows:
@@
The pseudo-$R^{2}$ is calculated using the formula:
@@
where SSE is the sum of squared errors and SST is the total sum of squares. It indicates the proportion of variation in the dependent variable that is explained by the model. A higher pseudo-$R^{2}$ suggests better explanatory power.
4.2 Interpretability methods
Despite the strong predictive performance of the BART model in both regression and classification tasks, it is often considered a “black-box" predictor due to its lack of transparency and interpretability chipman2010bart . Kapelner and Bleich kapelner2016bartmachine incorporated visualization tools such as partial dependence plots (PDPs) into the bartMachine framework. However, while PDPs provide insight into average marginal effects, they may fall short in capturing more complex or localized feature interactions.
To complement PDPs and provide a more nuanced interpretation of model behavior, this study also employs SHapley Additive exPlanations (SHAP). These tools examine feature contributions from different theoretical perspectives, thereby enhancing the depth and reliability of the model interpretation, especially for complex interactions and heterogeneous effects.
@@ 4.2.1 Partial Dependence Plot
This study employs the Partial Dependence Plot (PDP) from the bartMachine framework to examine how a specific predictor affects the response variable on average, while controlling for other predictors. The PDP is based on the formulation introduced by Friedman friedman2001greedy :
@@ 4.2.2 SHapley Additive exPlanations
SHAP (SHapley Additive exPlanations) is grounded in the game-theoretic concept of Shapley values, which were originally proposed for model explanation by Štrumbelj & Kononenko strumbelj2011general ; strumbelj2014explaining . However, the approach only gained widespread attention after Lundberg & Lee lundberg2017unified introduced SHAP as a unified framework for feature attribution. Their work not only rebranded the method but also extended its theoretical foundations, connecting Shapley values with other model-agnostic interpretability techniques such as LIME.
@@
$\phi_{i}=\sum_{S\subseteq F\setminus\{i\}}\frac{|S|!\cdot(|F|-|S|-1)!}{|F|!}[f(S\cup i)-f(S)]$
where $F$ is the set of all features; $S$ is the subset of features excluding $i$; $f(S)$ is the expected prediction when only features in $S$ are known.
In practice, for models with a large number of input variables, exhaustively enumerating all feature subsets S to compute Shapley values is computationally infeasible. To address this challenge, we employed the FastSHAP R package developed by Greenwell and Greenwell greenwell2020fastshap , which provides an efficient approximation of Shapley values for any supervised learning model. Rather than calculating exact Shapley values by iterating over all possible feature coalitions, FastSHAP employs a Monte Carlo-based approximation method proposed by Strumbel & Kononenko strumbelj2014explaining . The approximated outputs, commonly referred to as SHAP values, preserve the theoretical properties of Shapley values while enabling scalable interpretation of complex black-box models.
In our analytical framework, we first use Partial Dependence Plots (PDPs) from the bart package to explore the global structure and interaction patterns of the model. Then, we apply SHAP values to conduct detailed mechanism analysis. This combined approach helps provide a more comprehensive understanding of how input features affect the predictions and enhances the transparency and credibility of BART.
@@ 5.1.1 Model evaluation
To ensure a realistic evaluation of the model’s predictive performance and to avoid data leakage, we used data from 2000 to 2020 as the training set and data from 2021 to 2022 as the test set. Additionally, we applied 10-fold cross-validation to evaluate the model’s generalization ability. The BART model initially trained with default hyperparameters achieved a pseudo-R² of 0.989 and an RMSE of 0.1 on the training set. The 10-fold cross-validation yielded a pseudo-R² of 0.9586 and an RMSE of 0.1886. On the test set, the model achieved a pseudo-R² of 0.8493 and an RMSE of 0.2826. These results indicate that even with default settings, BART demonstrates strong fitting ability and successfully captures most of the underlying data patterns.
@@
However, the model’s predictive performance on unseen data was relatively weaker, and the residuals failed the normality test (p-value=0.00091<0.05), indicating potential model misspecification or non-Gaussian error distribution. To address this, we used the built-in grid search function in the bartMachine package to tune the model’s hyperparameters. The tuned model showed improved performance on the training set, with a pseudo-R² of 0.9955 and an RMSE of 0.06. Importantly, the cross-validated pseudo-R² remained stable at 0.9608, suggesting stable performance without signs of overfitting. Most notably, the pseudo-R² on the test set increased from 0.8493 to 0.8548, confirming that hyperparameter tuning enhanced the model’s predictive accuracy on new data. Moreover, the residuals from the tuned model passed the normality test (p-value = 0.269 >> 0.05), indicating a better fit to the data.
@@
In summary, the tuned BART model demonstrates both high predictive power and robust generalization. Therefore, all subsequent analyses on the relationships between green finance, climate risk, and carbon emission intensity are based on the tuned model.
5.1.2 Model Diagnostics and Statistical Significance Testing
@@
Figure 1 presents the residuals Q-Q plot to assess normality and the residuals vs predicted values plot to assess homoskedasticity. The Shapiro-Wilk test for residuals yielded a p-value of 0.269 >> 0.05, indicating that the residual distribution generally conforms to normality. Although a few points deviate in the tails of the Q-Q plot, the majority of points align closely along the diagonal line. Meanwhile, the residuals vs predicted values plot shows no obvious pattern or trend, suggesting that there is no significant heteroskedasticity issue.
For the overall model test, we permuted the response variable $y$, thereby breaking any association between the predictors and the outcome. This is equivalent to an omnibus F-test in linear regression. The result shows a p-value of 0, well below the conventional 0.05 threshold, indicating that the model’s predictive ability is highly statistically significant.
Also, we conducted individual significance tests for GFI and CPRI. In each case, the values of the target variable were randomly permuted within the sample, destroying its relationship with the response while keeping all other variables fixed. This allows us to test whether the variable contributes predictive power after controlling for others. The results show that GFI has a p-value of 0.0396 < 0.05, indicating a statistically significant effect on carbon emission intensity. In contrast, CPRI yields a p-value of 0.238, which exceeds the 0.05 threshold. Therefore, we fail to reject the null hypothesis and conclude that there is insufficient statistical evidence to support a significant effect of CPRI on carbon emission intensity.
Also, we conducted individual significance tests for GFI and CPRI. In each case, the values of the target variable were randomly permuted within the sample, destroying its relationship with the response while keeping all other variables fixed. This allows us to test whether the variable contributes predictive power after controlling for others. The results show that GFI has a p-value of 0.02, indicating a statistically significant effect on carbon emission intensity. In contrast, CPRI yields a p-value of 0.281, which exceeds the 0.05 threshold. Therefore, we fail to reject the null hypothesis and conclude that there is insufficient statistical evidence to support a significant effect of CPRI on carbon emission intensity.
@@ 5.2 Feature importance and interaction terms
This section utilizes the built-in functionalities of bartMachine to examine the importance of individual predictors and their interactions in BART, Additionally, SHAP methods are employed to visualize nonlinear relationships and interactions among variables. Figures 2, 3, and 4 present the variable importance and interaction results extracted from bartMachine, while Figures 5, and 6 show visualizations based on SHAP. As shown in Figure 2, TEC appears most frequently in the model’s decision paths and has relatively narrow 95% credible intervals (CI), indicating that it contributes the most to explaining variations in CEI, with a high degree of certainty. GFI ranks fifth, and its corresponding 95% CI is also relatively narrow, suggesting that green finance exerts a significant and stable influence within the model. In contrast, CPRI ranks the lowest, with a near-zero contribution, implying limited predictive relevance. The statistical significance of each variable are further validated in Section 5.1.2.
Regarding interaction effects, Figure 3 displays the top 10 variable pairs with the strongest interactions. Among them, GFI is involved in several interactions, including GFI × TEC, GFI × population, and GFI × govern. This suggests that the impact of green finance on CEI largely depends on its interaction with other key variables. Notably, the GFI × TEC interaction ranks fifth overall and is the most prominent among GFI-related pairs. Therefore, the subsequent analysis will focus on the interaction between GFI and TEC to explore the mechanism through which green finance influences the marginal effect of energy consumption on carbon emission intensity.
@@ 5.3 Nonlinear and interaction effects analysis
5.3.1 Nonlinear and regional effects of GFI
Figures 4 and 5 jointly explore the nonlinear and heterogeneous effects of GFI on CEI, using PDP from BART and SHAP visualizations. Figure 4 displays the PDP of GFI, illustrating the average marginal effect of GFI on CEI while holding all other variables constant. The results reveal a clear inverted U-shaped pattern: at low levels of GFI, the PDP curve rises, indicating that the initial development of green finance is associated with an increase in CEI. However, once GFI exceeds a certain threshold (around 0.3), the curve begins to decline, suggesting that further development of green finance helps reduce CEI. This implies that green finance has greater emission reduction potential in more advanced stages. The relatively narrow 95% confidence interval further confirms the stability and robustness of this nonlinear relationship.
The SHAP-based analysis in Figure 5 further reveals regional heterogeneity in the observed nonlinear relationship. In Eastern and Central China, the inverted U-shape is more pronounced, with most observations concentrated within the GFI range of 0.2-0.4. After GFI exceeds 0.3, SHAP values in both regions show a general downward trend. However, in Central China, most SHAP values remain positive, indicating that although GFI has reached a relatively high level, its emission reduction effect has not yet fully materialized, possibly due to the green finance sector still being in its expansion or investment phase. In contrast, in Eastern China, SHAP values decline more significantly after GFI exceeds 0.3, with some observations even turning negative. This may reflect a more mature green finance system in the region, which is beginning to effectively curb carbon intensity. In Western China, GFI levels are more widely distributed. When GFI is relatively low, SHAP values are mostly negative, suggesting that even in its early stages, green finance has already contributed to reducing CEI to some extent. As GFI increases, SHAP values gradually rise and approach positive territory, indicating that the emission reduction effect weakens during the mid-development stage. However, when GFI reaches approximately 0.3, SHAP values begin to decline again, showing signs of diminishing marginal returns or a “saturation effect.”
These findings confirm that the effectiveness of green finance in reducing carbon emissions is nonlinear and regionally differentiated, highlighting the importance of region-specific strategies in promoting green finance development.
@@ 5.3.1 Nonlinear and regional effects of GFI
Figure 4: PDP of GFI on CEI. The black line represents the average marginal effect of GFI on CEI when all other variables are held constant. The shaded gray area and blue lines indicate the 95% credible interval, reflecting model uncertainty. The narrowing of the credible interval suggests increasing model certainty. The x-axis shows GFI values at specified quantiles (5%, 10%, …, 95%), while the y-axis represents the average predicted CEI at each GFI level, conditional on other variables being fixed.
@@
Figure 5: SHAP PDP of GFI on CEI, color-coded by region. Each point represents an observation, with the x-axis showing the value of GFI and the y-axis showing the corresponding SHAP value, which reflects the marginal contribution of GFI to the model prediction of CEI. SHAP values represent the marginal contribution of a given feature (GFI) to the model’s prediction of CEI. A positive SHAP value indicates that GFI increases predicted CEI (i.e., contributes positively), while a negative value suggests a suppressing effect (i.e., contributes negatively). The color of the points denotes regional classification: Eastern (red), Central (purple), and Western (yellow).

### v2

@@ 3.3.3 Socioeconomic and environmental factors
In addition, carbon emissions are also influenced by a range of socioeconomic and environmental factors. Building on the existing literature from Section 2, this study incorporates several variables into the model to better capture the potential relationships among key predictors. These include the level of economic development, industrial structure, research and development (R&D) intensity, urbanization level, population density, foreign direct investment, degree of government intervention, environmental regulation intensity.
In light of empirical studies that have confirmed the significant regional heterogeneity in the impact of green finance on carbon emissions, this study further divides the provinces into eastern, central, and western regions based on the traditional regional classification method, and incorporates this classification as a categorical variable in the model.
@@ 3.4 Data sources
All data are obtained from official sources, including the China Statistical Yearbook, China Energy Statistical Yearbook, Regional Statistical Yearbooks, CSMAR database, Wind database, the Environmental Status Bulletin, and the China Deep Data database. Table 2 provides a detailed description of all variables and sources.
@@ 3.5 Data preprocessing
To reduce right-skewness in the distribution of variables, we applied logarithmic transformations to the dependent variable (CEI) and to the selected independent variables, namely economic development (GDP), population density (population), industrial structure (industry), and environmental regulation (environment), all of which exhibited a skewness value greater than 2. We then employed Generalized Variance Inflation Factors (GVIF) to conduct a preliminary analysis of potential multicollinearity among explanatory variables. A GVIF value close to 1 indicates little to no multicollinearity, while values equal to or greater than 5 or 10 suggest strong multicollinearity obrien2007caution . Similarly, we consider the adjusted generalized standard error inflation factor (aGSIF) for which values exceeding 2.2 or 3.2 are also considered indicative of significant multicollinearity shabangu2025monthly .
@@ 4.1 Model specification
Originally proposed by chipman2010bart , Bayesian Additive Regression Trees (BART) is a flexible, nonparametric Bayesian ensemble method that represents the response variable $y$ as the sum of $m$ regression trees:
@@
The update of the tree structure $T_{i}$ is conducted using a Metropolis–Hastings step. New tree proposals are generated through operations such as growing or pruning nodes, modifying split rules, or swapping decision nodes. The acceptance probability is computed by integrating out the terminal node parameters $M_{i}$, which is tractable due to the use of conjugate priors. Once the new tree is accepted, $M_{i}$ is updated by drawing from its full conditional posterior distribution, which follows a normal distribution. The noise variance $\sigma^{2}$ is updated at the end of each iteration by drawing from a full conditional inverse-gamma distribution. The MCMC algorithm was run for a large number of iterations, discarding the initial samples as burn-in. Satisfactory convergence was checked using trace plots and standard convergence diagnostics. From the resulting posterior samples, a sequence of fitted functions is obtained:
@@
The model fit is evaluated using Root Mean Square Error (RMSE) and pseudo-$R^{2}$. A lower RMSE indicates a better fit to the data. The RMSE is calculated as follows:
@@
whereas the pseudo-$R^{2}$ is calculated using the formula:
@@
Here, SSE is the sum of squared errors and SST is the total sum of squares, indicating the proportion of variation in the dependent variable that is explained by the model. A higher pseudo-$R^{2}$ suggests better explanatory power.
4.2 Summary tools
Despite the strong predictive performance of the BART model in both regression and classification tasks, As chipman2010bart mentioned, it is often considered a “black-box" predictor due to its complex model structure which cannot be easily summarized. kapelner2016bartmachine incorporated visualization tools such as partial dependence plots (PDPs) into the bartMachine framework. However, while PDPs provide insight into average marginal effects, they may fall short in capturing more complex or localized feature interactions.
To complement PDPs and provide a more nuanced interpretation of model behavior, this study also employs SHapley Additive exPlanations lundberg2017unified . These tools examine feature contributions from different perspectives, thereby enhancing the depth and reliability of the model interpretation, especially for complex interactions and heterogeneous effects.
@@ 4.2.1 Partial dependence plot
We adopt PDPs from the bartMachine framework to examine how a specific predictor affects the response variable on average, while controlling for other predictors. The PDP is based on the formulation introduced by friedman2001greedy :
@@
where $f$ represents the true but unknown model function, and $dP(x_{-j})$ denotes the marginal distribution of all covariates except the jthj-th one. The function $f_{j}$ measures the marginal or average effect of a single predictor $x_{j}$ on the response variable. Its principle is to “average out” the influence of the other variables by integrating over their marginal distribution. In other words, $f_{j}(x_{j})$ describes how the expected model prediction changes as $x_{j}$ varies, while the effects of the other variables are held at their average levels, thereby isolating the independent contribution of a single predictor to the model’s output.
@@ 4.2.2 SHapley Additive exPlanations
SHAP (SHapley Additive exPlanations) is a method grounded in the game-theoretic concept of Shapley values, which were originally proposed for model explanation by strumbelj2011general ; strumbelj2014explaining . However, the approach only gained widespread attention after lundberg2017unified introduced SHAP as a unified framework for feature attribution. Their work not only rebranded the method but also extended its theoretical foundations, connecting Shapley values with other model-agnostic interpretability techniques such as LIME ribeiro2016why .
@@
$\phi_{i}=\sum_{S\subseteq V\setminus\{i\}}\frac{|S|!\cdot(|V|-|S|-1)!}{|V|!}[f(S\cup i)-f(S)].$
where $V$ is the set of all features; $S$ is the subset of features excluding $i$; $f(S)$ is the expected prediction when only features in $S$ are known.
In practice, for models with a large number of input variables, exhaustively enumerating all feature subsets S to compute Shapley values is computationally infeasible. To address this challenge, we employed the FastSHAP R package developed by greenwell2020fastshap , which provides an efficient approximation of Shapley values for any supervised learning model. Rather than calculating exact Shapley values by iterating over all possible feature coalitions, FastSHAP employs the Monte Carlo-based approximation method proposed by strumbelj2014explaining . The approximated outputs, commonly referred to as SHAP values, preserve the theoretical properties of Shapley values while enabling scalable interpretation of complex black-box models.
In our analytical framework, we first use Partial Dependence Plots (PDPs) from the bart package to explore the global structure and interaction patterns of the model. Then, we complement PDPs with SHAP values to conduct a more in depth analysis of the interactions between variables analysis. This combined approach helps to provide a more comprehensive understanding of how input features affect the predictions, thus enhancing the transparency and reliability of BART.
@@ 5.1 Model performance
To ensure a realistic evaluation of the model’s predictive performance and to avoid data leakage, we used data from 2000 to 2020 as the training set and data from 2021 to 2022 as the test set. Additionally, we applied 10-fold cross-validation to evaluate the model’s generalization ability. The BART model initially trained with default hyperparameters achieved a pseudo-R² of 0.989 and an RMSE of 0.1 in the training set. The 10-fold cross-validation yielded a pseudo-R² of 0.9586 and an RMSE of 0.1886. In the test set, the model achieved a pseudo-R² of 0.8493 and an RMSE of 0.2826. These results indicate that even with default settings, BART achieves a good fit strong fitting ability and successfully captures most of the underlying data patterns.
@@
However, the model’s predictive performance on unseen data was relatively weaker, and the residuals failed the normality test (p-value=0.00091 << 0.05), indicating potential model misspecification or non-Gaussian error distribution. To address this, we used the built-in grid search function in the bartMachine package to tune the model’s hyperparameters. The tuned model showed improved performance on the training set, with a pseudo-R² of 0.9955 and an RMSE of 0.06. Importantly, the cross-validated pseudo-R² remained stable at 0.9608, suggesting stable performance without signs of overfitting. Most notably, the pseudo-R² on the test set increased from 0.8493 to 0.8548, confirming that hyperparameter tuning enhanced the model’s predictive accuracy on new data. Moreover, the residuals from the tuned model passed the normality test (p-value = 0.269 >> 0.05), indicating a better fit to the data.
@@
In summary, the tuned BART model seems to achieve both high predictive power and robust generalization. Therefore, all subsequent analyses on the relationships between green finance, climate risk, and carbon emission intensity are based on the tuned model.
5.1.1 Preliminary model diagnostics
@@
Figure 1 presents the residuals Q-Q plot to assess normality and the residuals vs predicted values plot to assess homoskedasticity. The Shapiro-Wilk test for residuals yielded a p-value of 0.269 >> 0.05, indicating that the assumption of normality should not be rejected. Although a few points deviate in the tails of the Q-Q plot, the majority of points align closely along the diagonal line. Meanwhile, the residuals vs predicted values plot shows no obvious pattern or trend, suggesting that there are no particular concerns regarding heteroskedasticity.
@@ 5.2 Feature importance and interaction terms
For the overall model test, we shuffled the order of the response variable $y$, thereby breaking any association between the predictors and the outcome. This is equivalent to an omnibus F-test in linear regression. The result shows a p-value of 0, well below the conventional 0.05 threshold, indicating that the model’s predictive ability is highly statistically significant.
Also, we conducted individual significance tests for GFI and CPRI. In each case, the values of the target variable were randomly permuted within the sample, destroying its relationship with the response while keeping all other variables fixed. This allows us to test whether the variable contributes predictive power after controlling for the others. The results show that GFI has a p-value of 0.0396 << 0.05, indicating a statistically significant effect on carbon emission intensity. In contrast, CPRI yields a p-value of 0.238, which exceeds the 0.05 threshold. Therefore, we fail to reject the null hypothesis and conclude that there is insufficient statistical evidence to support a significant effect of CPRI on carbon emission intensity.
5.3 Feature importance and interaction terms
This section utilizes the built-in functionalities of bartMachine to examine the importance of individual predictors and their interactions within BART. In the BART framework, variable importance is evaluated using the inclusion proportion metric chipman2010bart ; bleich2014variable . The inclusion proportion for a given predictor represents the frequency with which that variable is selected as a splitting rule across all posterior samples (among all tree structures generated during the MCMC iterations). A higher inclusion proportion indicates that the variable appears more frequently in the model’s partitioning process and thus contributes more substantially to explaining variations in the response variable.
Interaction effects are evaluated following the approach of damien2013bayesian , where an interaction between two predictors is considered present when both variables appear within splitting rules along a path in the same tree. To quantify these interactions, the method counts how often each pair of variables co-occurs in such splits across all trees and posterior samples. Summing these counts provides an overall measure of the interaction strength for each pair, which can then be used to assess the relative importance of their combined effect on the response. This approach allows the identification of predictor pairs whose joint influence is substantial, including cases where a variable may interact with itself.
@@ 5.3 Feature importance and interaction terms
As shown in Figure 2, TEC appears most frequently in the model’s decision paths and has relatively narrow 95% credible intervals, indicating that it contributes the most to explaining variations in CEI, with a high degree of certainty. GFI ranks fifth, and its 95% credible interval is relatively wide compared with those of other predictors, suggesting that although green finance exerts a significant influence, the model exhibits some uncertainty about the precise magnitude of its effect. In contrast, CPRI ranks the lowest, with a near-zero inclusion proportion and one of the narrowest 95% credible interval, suggesting that it plays a limited role in explaining variations in CEI and that the model is fairly certain about this result. The statistical significance of each variable are further validated in Section 5.1.1.
In terms of interaction effects, bartMachine constructs 25 independent BART models, each trained from a different random initialization (randomly chosen initial tree structure and parameter state). The interaction strengths are then averaged across these models to obtain stable estimates that account for different posterior modes in the tree and distribution. Figure 3 displays the top 10 variable pairs with the strongest interactions, as measured by their average co-occurrence across decision paths in 25 constructed models. Among these, GFI appears frequently as a component of key interacting pairs, including GFI × TEC, GFI × population, and GFI × govern. The presence of GFI in multiple high-ranking interactions indicates that its influence on CEI may be highly conditional on other structural and socioeconomic factors. In particular, the interaction between GFI and TEC ranks fifth overall and stands out as the most influential GFI-related pair. This finding suggests that the marginal impact of green finance on carbon emissions may be amplified or moderated depending on the level of energy consumption. Therefore, the subsequent analysis will focus on the interaction between GFI and TEC to explore the mechanism through which green finance influences the marginal effect of energy consumption on carbon emission intensity.
@@ 5.4 Nonlinear and interaction effects analysis
Figures 4 and 5 jointly explore the nonlinear and heterogeneous effects of GFI on CEI, using PDP from BART and SHAP visualizations. The PDP curve in Figure 4 exhibits a typical inverted U-shaped pattern: at lower GFI levels, CEI increases with GFI, but once GFI exceeds a threshold of approximately 0.3, the curve begins to decline, indicating an emission-reduction effect at higher levels of GFI. However, the interpretability of this nonlinear pattern depends on the underlying data-generating process.
@@
Figure 4: PDP of GFI on CEI. The black line represents the average marginal effect of GFI on CEI when all other variables are held constant. The shaded gray area and blue lines indicate the 95% credible interval, reflecting model uncertainty. The narrowing of the credible interval suggests increasing model certainty. The x-axis shows GFI values, while the y-axis represents the average predicted CEI at each GFI level, conditional on other variables being fixed.
@@
Figure 5: SHAP PDP of GFI on CEI, color-coded by region (red for Eastern, purple for Central, yellow for Western). Each point represents an observation, with the x-axis showing the value of GFI and the y-axis showing the corresponding SHAP value, which reflects the marginal contribution of GFI to the model prediction of CEI. SHAP values represent the marginal contribution of a given feature (GFI) to the model’s prediction of CEI. A positive SHAP value indicates that GFI increases predicted CEI (i.e., contributes positively), while a negative value suggests a suppressing effect (i.e., contributes negatively).
The PDP is constructed by averaging over the effects of other variables while allowing GFI to vary, which implicitly assumes that GFI is independent or only weakly correlated with other covariates. In real economic systems, this assumption rarely holds, particularly in the low-GFI range where green finance is often closely related to factors such as technological progress, industrial structure, and policy implementation. Therefore, the apparent “positive correlation” observed at low GFI levels is more likely to reflect unobserved confounding effects rather than the true marginal impact of GFI. For example, in regions with lower technological development or less optimized economic structures, both insufficient green finance and higher carbon emissions tend to occur simultaneously, resulting in a spurious positive relationship in the model.
Once GFI exceeds the threshold of approximately 0.3, the influence of these confounding factors gradually weakens, and the independent role of green finance begins to emerge. The downward trend in this range more likely reflects the genuine emission-reduction effect of green finance. On one hand, a higher level of GFI channels financial and policy resources toward low-carbon, energy-efficient, and renewable energy projects, directly lowering carbon emission intensity. On the other hand, the expansion of green finance may also stimulate technological innovation and green industrial upgrading, establishing a long-term mitigation mechanism. In addition, a mature green finance system can enhance the environmental governance capacity of enterprises and local governments, leading to more efficient energy use and emission control. Therefore, the most meaningful part of Figure 4 lies in the high-GFI range, where green finance demonstrates a tangible and economically interpretable marginal effect on reducing CEI. This indicates that the environmental benefits of green finance become effective only when it reaches a sufficiently advanced level of development.
Figure 5 presents the SHAP-based analysis, which further reveals the regional heterogeneity underlying the nonlinear relationship observed in the PDP results. As shown in the figure, most samples in the range of GFI below approximately 0.3 are concentrated in the central and western regions, whereas those from the eastern region are primarily distributed at higher GFI levels above 0.3. It is important to note that although the SHAP values for the western region are generally negative at low GFI levels, this does not imply that low GFI exerts a genuine emission-reduction effect. As discussed in the above, low GFI levels are often accompanied by limited technological innovation capacity, slower industrial upgrading, and weaker policy implementation. Therefore, the relationship observed in this range likely reflects unobserved structural factors or confounding effects rather than the true marginal contribution of GFI.
Beyond the threshold of approximately 0.3, the carbon reduction effect of GFI becomes evident and displays notable regional heterogeneity. In the central and western regions, SHAP values exhibit an overall downward trend but remain mostly positive, suggesting that although green finance has developed to a higher level, its carbon reduction potential has not yet been fully realized. These regions may still be in a stage dominated by investment expansion and institutional development, where economic foundations, technological capabilities, and policy frameworks require further improvement. In contrast, SHAP values in the eastern region show a more pronounced decline at higher GFI levels, with some observations even turning negative. This indicates that in regions with stronger economic foundations, greater technological innovation, and more mature policy systems, advanced green finance development has begun to exert a substantial and measurable mitigating effect on carbon emission intensity, reflecting the effectiveness of a well-established and efficient green finance system.
Taken together with the PDP results, these findings demonstrate that the carbon reduction effects of green finance are not only nonlinear but also regionally heterogeneous. More importantly, the genuine marginal impact of green finance tends to emerge only once it surpasses a certain development threshold, and both the strength and timing of this effect vary significantly across regions. This evidence underscores the need for region-specific strategies that align with local economic structures and development stages to promote green finance as a sustained mechanism for reducing carbon intensity.